OpenStudy (lil.bossy):

which combination of integers can be used to generate the Pythagorean triple (8,15,17)? Please can someone help me. i don't know how to solve this.

jhonyy9 (jhonyy9):

do you know what mean Pythagorean triple ?

OpenStudy (lil.bossy):

No I do not

jhonyy9 (jhonyy9):

collaborate please

jhonyy9 (jhonyy9):

from geometry in case of a right triangle do you know what say the pythagora's theorem

jhonyy9 (jhonyy9):

?

OpenStudy (lil.bossy):

what would the equation I use be? my choices are x and y values and the pythagorean theorem is a^2+b^2=c^2

jhonyy9 (jhonyy9):

the first pythagorean triple are (3,4,5) and we know that from these can be generated pythegorean triple too do you know it how ?

OpenStudy (lil.bossy):

No i do not will you please explain

jhonyy9 (jhonyy9):

from these triplet of (3,4,5) result that every (3n,4n,5n) will be too pythagorean triplet

OpenStudy (lil.bossy):

but where do the x and y's come in

jhonyy9 (jhonyy9):

but than you check these (8,15,17) not will come never from these triplet of (3,4,5)

jhonyy9 (jhonyy9):

@phi @Zarkon please any idea ?

OpenStudy (phi):

I am guessing they are referring to https://en.wikipedia.org/wiki/Formulas_for_generating_Pythagorean_triples#Generating_triples_using_quadratic_equations and in this case, we use 2 m n = 8 m n = 4 so either 1,4 or 2,2 is the pair they are looking for

jhonyy9 (jhonyy9):

8^2 +15^2 =17^2 how we can generating these triplet ?

OpenStudy (phi):

using m=1, n= 4 the other two sides are n^2 - m^2 = 16-1 = 15 and n^2+m^2 = 16+1 = 17 and the first side is 2mn= 2*1*4= 8 so (1,4) looks like the pair we want

jhonyy9 (jhonyy9):

@lil.bossy do you understand it now ? thank you @phi

OpenStudy (lil.bossy):

yes thank both so much